# Cross-Dataset Gaze Estimation by Evidential Inter-intra Fusion

### v1

@@ 3.2. Cross-dataset Branch
To improve performance in unseen domains, preliminary experiments have demonstrated that utilizing data from various source domains can significantly boost the model’s effectiveness in such scenarios. This insight leads to the creation of the cross-dataset branch, a key component designed to combine information from different data sources. The cross-dataset branch is equipped with multiple Multi-Feature Fusion (MFF) modules and regression layers, whose architecture mostly mirrors the higher layers of the single-dataset branch. The intuition behind this design is that the lower layers usually capture generalizable features across different datasets and higher layers capture dataset-specific features, so we preserve the features from lower layers, and only fuse the features from high layers for tuning the cross-dataset branch.
@@
Formally, let $f_{n}^{l}$ denote the $l$-th layer feature of the $n$-th source dataset branch. Assuming fusion of features from multiple source datasets begins at the $K$-th layer, the fusion process of the MFF module can be represented as:
(3) $f_{\text{cross}}^{K}=\text{MFF}_{k}(f_{1}^{K},f_{2}^{K},\ldots,f_{S_{N}}^{K}),$
@@
(4) $f_{\text{cross}}^{k+1}=\text{MFF}_{k}(f_{1}^{k},f_{2}^{k},\ldots,f_{S_{N}}^{k},f_{\text{cross}}^{k}),$
@@
The inputs of MFF module come from different branches at the same layer, so their features are spatially aligned. But they may not be aligned in channels, hence, we first independently apply multiple point-wise convolutions to align features in channels. Subsequently, we use the Attention Feature Fusion (AFF) module proposed by Dai et al. (Dai et al., 2021) to fuse features pairwise (refer to (Dai et al., 2021) for details). The computation of $k$-th MFF module is expressed as:
(5) $\text{MFF}_{k}=\text{AFF}(...,\text{AFF}(\text{conv}(\cdot),\text{conv}(\cdot)),...,\text{conv}(\cdot)).$
3.3. Evidential Fusion Module
In this section, we introduce both an intra-evidential fusion module for fusing multiple local regressors within each dataset and an inter-evidential fusion module for fusing multiple dataset branches. The evidential fusion module is based on evidence regression learning (Amini et al., 2020). From its perspective, the gaze component $y$ is sampled from a normal distribution, with the mean and variance sampled from normal and Inverse-Gamma (NIG) distributions, respectively. Therefore, $y$ can be seen as indirectly sampled from an NIG distribution with parameters $\boldsymbol{\tau}=(\delta,\gamma,\alpha,\beta)$, where $\delta\in\mathbb{R}$, $\gamma>0$, $\alpha>1$, and $\beta>0$. This can be expressed as:
(6) $y\sim\mathcal{N}(\mu,\sigma^{2}),\quad\mu\sim\mathcal{N}(\delta,\sigma^{2}\gamma^{-1}),\quad\sigma^{2}\sim\Gamma^{-1}(\alpha,\beta),$
where $\Gamma(\cdot)$ is a gamma function. All of our local regressors adopt the evidential approach which is implemented by a fully connected layer that outputs four numbers corresponding to $\delta$, $\gamma$, $\alpha$, and $\beta$. During training, the evidence learning loss $\mathcal{L}_{\text{evidence}}$ consists of two essential components: the negative log-likelihood $\mathcal{L}_{\text{NLL}}$ and a regularization term $\mathcal{L}_{\text{R}}$. The negative log-likelihood $\mathcal{L}_{\text{NLL}}$ is defined as:
(7) $\begin{split}\mathcal{L}_{\text{NLL}}(\delta,\gamma,\alpha,\beta)&=\frac{1}{2}\log\left(\frac{\pi}{\gamma}\right)-\alpha\log\left(\Omega\right)\\ &\quad+\left(\alpha+\frac{1}{2}\right)\log\left((y-\delta)^{2}\gamma+\Omega\right)\\ &\quad+\log\left(\frac{\Gamma(\alpha)}{\Gamma\left(\alpha+\frac{1}{2}\right)}\right),\end{split}$
@@ 3.3. Evidential Fusion Module
(8) $\mathcal{L}_{\text{R}}(\delta,\gamma,\alpha,\beta)=|y-\delta|\cdot(2\gamma+\alpha).$
@@
(9) $\mathcal{L}_{\text{evidence}}(\delta,\gamma,\alpha,\beta)=\mathcal{L}_{\text{NLL}}(\delta,\gamma,\alpha,\beta)+\lambda\mathcal{L}_{\text{R}}(\delta,\gamma,\alpha,\beta),$
@@
Building upon evidence regression learning, the evidential fusion module employs the Mixture of normal-inverse gamma distribution (MoNIG) (Ma et al., 2021) for fusion. The resulting distribution remains NIG, represented as:
(10) $\begin{split}\text{MoNIG}(\delta,\gamma,\alpha,\beta)=&\text{NIG}(\delta_{1},\gamma_{1},\alpha_{1},\beta_{1})\oplus\\ &\text{NIG}(\delta_{2},\gamma_{2},\alpha_{2},\beta_{2})\oplus\\ &\cdots\oplus\text{NIG}(\delta_{M},\gamma_{M},\alpha_{M},\beta_{M}),\end{split}$
@@
(11) $\begin{split}\delta&=\frac{\sum_{i=1}^{M}\gamma_{i}\delta_{i}}{\sum_{i=1}^{M}\gamma_{i}},\quad\gamma=\sum_{i=1}^{M}\gamma_{i},\\ \alpha&=\sum_{i=1}^{M}\alpha_{i}+\frac{1}{M},\quad\beta=\sum_{i=1}^{M}\beta_{i}+\frac{1}{M}\sum_{i=1}^{M}\gamma_{i}(\delta_{i}-\delta)^{2}.\end{split}$
@@
(12) $\begin{split}\text{NIG}_{\text{BR}}^{n}(\delta_{\text{BR}}^{n},\gamma_{\text{BR}}^{n},\alpha_{\text{BR}}^{n},\beta_{\text{BR}}^{n})=&\text{NIG}_{\text{LR}_{1}}^{n}\oplus\\ &\text{NIG}_{\text{LR}_{2}}^{n}\oplus\cdots\ \oplus\text{NIG}_{\text{LR}_{G}}^{n}.\end{split}$
@@
(13) $\begin{split}\text{NIG}(\delta,\gamma,\alpha,\beta)=&\text{NIG}_{\text{BR}}^{1}\oplus\ldots\\ &\oplus\text{NIG}_{\text{BR}}^{S_{N}}\oplus\text{NIG}_{\text{BR}}^{cross},\end{split}$
@@
(14) $\underbrace{\mathbb{E}[\mu]=\delta}_{\text{gaze}},\qquad\underbrace{\mathbb{E}[\sigma^{2}]=\tfrac{\beta}{\alpha-1}}_{\text{aleatoric}},\qquad\underbrace{\text{Var}[\mu]=\tfrac{\beta}{\gamma(\alpha-1)}}_{\text{epistemic}}.$
3.4. Two-stage Training Process
To enable the model to rapidly adapt to variations in the number of datasets, we devise a two-stage training strategy. In stage 1, we train each source domain branch separately. For each dataset branch, the corresponding source domain data is divided into $G$ gaze groups $\mathcal{I}^{g}$ based on Equation 2. The output of $G$ local regressors for the $s$-th sample of the source domain data is denoted as $\delta_{\text{LR}_{g}}^{s},\gamma_{\text{LR}_{g}}^{s},\alpha_{\text{LR}_{g}}^{s},\beta_{\text{LR}_{g}}^{s}$. We only compute $\mathcal{L}_{\text{local}}^{g}$ within the $g$-th gaze group data subspace corresponding to the $g$-th local regressor, as follows:
(15) $\mathcal{L}_{\text{local}}^{g}=\sum_{y^{s}\in\mathcal{I}^{g}}\mathcal{L}_{\text{evidence}}(\delta_{\text{LR}_{g}}^{s},\gamma_{\text{LR}_{g}}^{s},\alpha_{\text{LR}_{g}}^{s},\beta_{\text{LR}_{g}}^{s},y^{s}).$
Here, $y^{s}$ represents the gaze label of the $s$-th sample. Then, we compute the output of this dataset branch after intra-evidential fusion according to Equation 12, denoted as $\delta_{\text{BR}}^{s},\gamma_{\text{BR}}^{s},\alpha_{\text{BR}}^{s},\beta_{\text{BR}}^{s}$. We compute the loss $\mathcal{L}_{\text{global}}$ for it across the entire source domain dataset as follows:
(16) $\mathcal{L}_{\text{global}}=\sum\mathcal{L}_{\text{evidence}}(\delta_{\text{BR}}^{s},\gamma_{\text{BR}}^{s},\alpha_{\text{BR}}^{s},\beta_{\text{BR}}^{s},y^{s})$
@@ 3.4. Two-stage Training Process
(17) $\mathcal{L}_{\text{branch}}=\sum_{g=1}^{G}\mathcal{L}_{\text{local}}^{g}+\mathcal{L}_{\text{global}}.$
@@
In stage 2, we create a cross-dataset branch and jointly train all branches for only a few epochs. Specifically, we first merge multiple source domain datasets and then use Equation 2 to obtain $G$ gaze groups $\mathcal{I}^{g}_{\text{cross}}$ for the cross-dataset branch. Then, we compute the loss $\mathcal{L}_{\text{cross}}$ for the cross-dataset branch according to Equation 17. Given the outputs of each single branch, we fuse them using Equation 13 to obtain $\delta^{s},\gamma^{s},\alpha^{s},\beta^{s}$, and then compute the loss $\mathcal{L}_{\text{joint}}$ across all source domain datasets as follows:
@@
(18) $\mathcal{L}_{\text{joint}}=\sum_{\mathcal{D}_{S_{1}}...\mathcal{D}_{S_{N}}}\mathcal{L}_{\text{evidence}}(\delta^{s},\gamma^{s},\alpha^{s},\beta^{s},y^{s}),$
Therefore, the total loss for stage 2 is
(19) $\mathcal{L}_{\text{stage2}}=\mathcal{L}_{\text{cross}}+\mathcal{L}_{\text{joint}}.$
4. Experiment
4.1. Cross-dataset Setting
@@
The experiments utilize four widely used gaze estimation datasets, EyeDiap (Funes Mora et al., 2014), MPIIFaceGaze (Zhang et al., 2017), Gaze360 (Kellnhofer et al., 2019), and ETH-XGaze (Zhang et al., 2020). We pre-process the data following the technique in (Cheng et al., 2021). To facilitate comparisons, we reference the setup of cross-domain scenarios (Xu et al., 2023; Cai et al., 2023) in our cross-dataset setting, i.e., we choose ETH-XGaze and Gaze360, known for their wider gaze distributions, as the source domains, and MPIIGaze and EyeDiap as the unseen datasets. Specific details regarding the datasets are provided below.
ETH-Gaze (Source Domain): This dataset is captured in a laboratory environment using high-resolution cameras, comprising 756k training images from 80 subjects. Since the ETH-Gaze dataset lacks gaze target annotations for the test set, we redefine the dataset splits following the setup in (Cheng et al., 2021). Specifically, we split the original training set into two subsets and use them as training and testing sets, as done in previous works (Cheng and Lu, 2023).
Gaze360 (Source Domain): This dataset is collected in both indoor and outdoor environments using a 360-degree camera. The dataset contains 84k training images from 54 subjects and 16k testing images from 15 subjects.
MPIIFaceGaze (Unseen Domain): This dataset is collected during natural laptop use, with images captured using the laptop camera. It comprises 45k images from 15 users.
EyeDiap (Unseen Domain): This dataset is collected in a laboratory environment, where participants are asked to gaze at two types of visual target sessions: screen target and 3D floating ball. It comprises a total of 16k images gathered from 16 users.
4.2. Implementation Details
@@
We use ResNet-18 backbone following previous studies (Cheng et al., 2022; Xu et al., 2023; Cai et al., 2023; Bao and Lu, 2023). All images are resized to 224 × 224 . The batch size is 128, and we use the Adam optimizer with a learning rate of $10^{-4}$. In stage 1, we train on each source domain dataset for 100k batches. In stage 2, we perform joint training for only 5k batches using all source domain datasets. Considering the discrepancy in the quantity of data among source domain datasets, during the joint training in stage 2, instead of directly mixing multiple datasets for sampling, we adopt a balanced mixing approach, where the datasets with fewer samples are oversampling to balance the data across all source domain datasets. By default, we set the number of local regressors $G$ to 8, the overlap coefficient $\alpha$ to 2.0, the evidence learning regularization loss $\lambda$ to 0.01, and the starting MFF layer $K$ in the cross-dataset branch to 3.
4.3. Domain Generalization Results
We trained our method on the source domains (ETH-Gaze and Gaze360 datasets) and compared it with three categories of methods:
@@ 4.3. Domain Generalization Results
Table 1. Comparison of gaze estimation performance in terms of angle error (°) in domain generalization setting. The best, second best and third best results are denoted as red, blue and green respectively.
Method Test Source Domain Unseen Domain Average
Gaze360 ETH-Gaze MPIIFaceGaze EyeDiap Unseen All
Baseline (ETH-Gaze) 19.98 5.02 8.59 7.87 8.23 10.37
Baseline (Gaze360) 11.42 17.43 7.45 10.36 8.91 11.67
Baseline (ETH-Gaze+Gaze360) 11.45 5.11 7.01 7.39 7.20 7.74
PureGaze (ETH-Gaze) - - 7.08 7.48 7.28 -
PureGaze (Gaze360) - - 9.28 9.32 9.30 -
GCF (ETH-Gaze) - - 6.50 7.44 6.97 -
GCF (Gaze360) - - 7.55 9.03 8.29 -
EIF (ETH-Gaze+Gaze360) 11.03 4.75 6.40 6.77 6.58 7.24
Single-dataset baseline (ETH-Gaze, Gaze360):

### v2

@@ 3.2 Cross-dataset Branch
To improve performance in target domains, preliminary experiments have demonstrated that utilizing data from various source domains can significantly boost the model’s effectiveness in such scenarios. This insight leads to the creation of the cross-dataset branch, a key component designed to combine information from different data sources. The cross-dataset branch is equipped with multiple Multi-Feature Fusion (MFF) modules and regression layers, whose architecture mostly mirrors the higher layers of the single-dataset branch.
@@
Formally, let $f_{n}^{l}$ denote the $l$-th layer feature of the $n$-th source dataset branch. Assuming the feature fusion from multiple source datasets begins at the $K$-th layer, the fusion process of the MFF module can be represented as:
$f_{\text{cross}}^{K}=\text{MFF}_{k}(f_{1}^{K},f_{2}^{K},\ldots,f_{S_{N}}^{K}),$ (4)
@@
$f_{\text{cross}}^{k+1}=\text{MFF}_{k}(f_{1}^{k},f_{2}^{k},\ldots,f_{S_{N}}^{k},f_{\text{cross}}^{k}),$ (5)
@@
The inputs of MFF module come from different branches at the same layer, so their features are spatially aligned. But they may not be aligned in channels, hence, we first independently apply multiple point-wise convolutions to align features in channels. Subsequently, we use the Attention Feature Fusion (AFF) module proposed by Dai et al. [12] to fuse features pairwise (refer to [12] for details). The computation of $k$-th MFF module is expressed as:
$\text{MFF}_{k}=\text{AFF}(...,\text{AFF}(\text{conv}(\cdot),\text{conv}(\cdot))).$ (6)
3.3 Evidential Fusion Module
In this section, we introduce both an intra-evidential fusion module for fusing multiple local regressors within each dataset and an inter-evidential fusion module for fusing multiple dataset branches. The evidential fusion module is based on evidence regression learning [2]. From its perspective, the gaze component $y$ is sampled from a normal distribution, with the mean and variance sampled from normal and Inverse-Gamma (NIG) distribution, respectively. Therefore, $y$ can be seen as indirectly sampled from an NIG distribution with parameters $\boldsymbol{\tau}=(\delta,\gamma,\alpha,\beta)$, where $\delta\in\mathbb{R}$, $\gamma>0$, $\alpha>1$, and $\beta>0$. This can be expressed as:
$y\sim\mathcal{N}(\mu,\sigma^{2}),\ \mu\sim\mathcal{N}(\delta,\sigma^{2}\gamma^{-1}),\ \sigma^{2}\sim\Gamma^{-1}(\alpha,\beta).$ (7)
where $\Gamma(\cdot)$ is a gamma function. All of our local regressors adopt the evidential approach which is implemented by a fully connected layer that outputs four numbers corresponding to $\delta$, $\gamma$, $\alpha$, and $\beta$. During training, the evidence learning loss $\mathcal{L}_{\text{evidence}}$ consists of two essential components: a negative log-likelihood $\mathcal{L}_{\text{NLL}}$ and a regularization term $\mathcal{L}_{\text{R}}$. The negative log-likelihood $\mathcal{L}_{\text{NLL}}$ is defined as:
$\begin{split}\mathcal{L}_{\text{NLL}}(\delta,\gamma,\alpha,\beta)&=\frac{1}{2}\log\left(\frac{\pi}{\gamma}\right)-\alpha\log\left(\Omega\right)\\ &\quad+\left(\alpha+\frac{1}{2}\right)\log\left((y-\delta)^{2}\gamma+\Omega\right)\\ &\quad+\log\left(\frac{\Gamma(\alpha)}{\Gamma\left(\alpha+\frac{1}{2}\right)}\right),\end{split}$ (8)
@@ 3.3 Evidential Fusion Module
$\mathcal{L}_{\text{R}}(\delta,\gamma,\alpha,\beta)=|y-\delta|\cdot(2\gamma+\alpha).$ (9)
@@
$\begin{split}\mathcal{L}_{\text{evidence}}(\delta,\gamma,\alpha,\beta)=&\mathcal{L}_{\text{NLL}}(\delta,\gamma,\alpha,\beta)\\ &+\lambda\mathcal{L}_{\text{R}}(\delta,\gamma,\alpha,\beta),\end{split}$ (10)
@@
Building upon evidence regression learning, the evidential fusion module employs the Mixture of normal-inverse gamma distribution (MoNIG) [26] for fusion. The resulting distribution remains NIG, represented as:
$\begin{split}\text{MoNIG}(\delta,\gamma,\alpha,\beta)=&\text{NIG}(\delta_{1},\gamma_{1},\alpha_{1},\beta_{1})\oplus\\ &\text{NIG}(\delta_{2},\gamma_{2},\alpha_{2},\beta_{2})\oplus\\ &\cdots\oplus\text{NIG}(\delta_{M},\gamma_{M},\alpha_{M},\beta_{M}),\end{split}$ (11)
@@
$\begin{split}\delta&=\frac{\sum_{i=1}^{M}\gamma_{i}\delta_{i}}{\sum_{i=1}^{M}\gamma_{i}},\quad\gamma=\sum_{i=1}^{M}\gamma_{i},\quad\alpha=\sum_{i=1}^{M}\alpha_{i}+\frac{1}{M},\\ \beta&=\sum_{i=1}^{M}\beta_{i}+\frac{1}{M}\sum_{i=1}^{M}\gamma_{i}(\delta_{i}-\delta)^{2}.\end{split}$ (12)
@@
$\begin{split}\text{NIG}_{\text{BR}}^{n}(\delta_{\text{BR}}^{n},\gamma_{\text{BR}}^{n},\alpha_{\text{BR}}^{n},\beta_{\text{BR}}^{n})=&\text{NIG}_{\text{LR}_{1}}^{n}\oplus\cdots\ \\ &\oplus\text{NIG}_{\text{LR}_{G}}^{n}.\end{split}$ (13)
@@
$\begin{split}\text{NIG}(\delta,\gamma,\alpha,\beta)=&\text{NIG}_{\text{BR}}^{1}\oplus\ldots\\ &\oplus\text{NIG}_{\text{BR}}^{S_{N}}\oplus\text{NIG}_{\text{BR}}^{cross},\end{split}$ (14)
@@
$\underbrace{\mathbb{E}[\mu]=\delta}_{\text{gaze}},\quad\underbrace{\mathbb{E}[\sigma^{2}]=\tfrac{\beta}{\alpha-1}}_{\text{aleatoric}},\quad\underbrace{\text{Var}[\mu]=\tfrac{\beta}{\gamma(\alpha-1)}}_{\text{epistemic}}.$ (15)
3.4 Two-stage Training Process
To enable the model to rapidly adapt to variations in the number of datasets, we devise a two-stage training strategy. In stage 1, we train each source domain branch separately. For each dataset branch, the corresponding source domain data is divided into $G$ gaze groups $\mathcal{I}^{g}$ based on Eq. 3. The output of $G$ local regressors for the $s$-th sample of the source domain data is denoted as $\delta_{\text{LR}_{g}}^{s},\gamma_{\text{LR}_{g}}^{s},\alpha_{\text{LR}_{g}}^{s},\beta_{\text{LR}_{g}}^{s}$. We only compute $\mathcal{L}_{\text{local}}^{g}$ within the $g$-th gaze group data subspace corresponding to the $g$-th local regressor, as follows:
$\mathcal{L}_{\text{local}}^{g}=\sum_{y^{s}\in\mathcal{I}^{g}}\mathcal{L}_{\text{evidence}}(\delta_{\text{LR}_{g}}^{s},\gamma_{\text{LR}_{g}}^{s},\alpha_{\text{LR}_{g}}^{s},\beta_{\text{LR}_{g}}^{s},y^{s}).$ (16)
Here, $y^{s}$ represents the gaze label of the $s$-th sample. Then, we compute the output of this dataset branch after intra-evidential fusion according to Eq. 13, denoted as $\delta_{\text{BR}}^{s},\gamma_{\text{BR}}^{s},\alpha_{\text{BR}}^{s},\beta_{\text{BR}}^{s}$. We compute the loss $\mathcal{L}_{\text{global}}$ for it across the entire source domain dataset as follows:
$\mathcal{L}_{\text{global}}=\sum\mathcal{L}_{\text{evidence}}(\delta_{\text{BR}}^{s},\gamma_{\text{BR}}^{s},\alpha_{\text{BR}}^{s},\beta_{\text{BR}}^{s},y^{s})$ (17)
@@ 3.4 Two-stage Training Process
$\mathcal{L}_{\text{branch}}=\sum_{g=1}^{G}\mathcal{L}_{\text{local}}^{g}+\mathcal{L}_{\text{global}}.$ (18)
@@
In Stage 2, we create a cross-dataset branch and jointly train all branches for a few epochs. We merge multiple source domain datasets and use Eq. 3 to obtain $G$ gaze groups, each characterized by center $\zeta_{\text{center}}^{g}$ and interval length $L^{g}$. The cross-dataset branch loss $\mathcal{L}_{\text{cross}}$ is computed according to Eq. 18, but utilizing all source datasets. Given the outputs of each branch, we fuse them using Eq. 14 to obtain $\delta,\gamma,\alpha,\beta$ and compute the joint loss $\mathcal{L}_{\text{joint}}$ across all source domain datasets:
@@
$\mathcal{L}_{\text{joint}}=\sum_{\mathcal{D}_{S_{1}}...\mathcal{D}_{S_{N}}}\mathcal{L}_{\text{evidence}}(\delta,\gamma,\alpha,\beta,y).$ (19)
Thus, the total loss for Stage 2 is:
$\mathcal{L}_{\text{stage2}}=\mathcal{L}_{\text{cross}}+\mathcal{L}_{\text{joint}}.$ (20)
In this approach, each single-dataset branch is trained independently in Stage 1, allowing for model reuse. Thus, various model combinations can be achieved by training only in Stage 2. Furthermore, to promote effective learning in the cross-dataset branch, each source domain dataset is split in a specific ratio across the two stages, ensuring that the data used in joint training remains distinct from that in Stage 1, thereby reducing the risk of overfitting and enhancing cross-dataset branch learning.
4 Experiment
4.1 Cross-dataset Setting
@@
The experiments utilize four widely used gaze estimation datasets, EyeDiap [14], MPIIFaceGaze [40], Gaze360 [20], and ETH-XGaze [41]. We pre-process the data following the technique in [10]. To facilitate comparisons, we reference the setup of cross-domain scenarios [36, 6] in our cross-dataset setting, i.e., we choose ETH-XGaze and Gaze360, known for their wider gaze distributions and larger data volumes, as the source domains, and MPIIGaze and EyeDiap as the target datasets. Detailed information about the datasets can be found in the Supplementary Material.
4.2 Implementation Details
@@
We use ResNet-18 backbone following previous studies [11, 6, 4]. All images are resized to 224 × 224, and we use the Adam optimizer with a learning rate of $10^{-4}$. In Stage 1, we train on each source domain dataset for 100k batches. Specifically, for the Gaze360 single-dataset branch, we use the ETH-Gaze’s backbone for pre-training, similar to the previous work [8]. In stage 2, we perform joint training for only 5k batches using all source domain datasets. Considering the discrepancy in the quantity of data among source domain datasets, during the joint training in stage 2, instead of directly mixing multiple datasets for sampling, we adopt a balanced mixing approach, where the datasets with fewer samples are oversampling to balance the data across all source domain datasets. Notably, our two-stage learning strategy allows for quick adaptation. Once the single-dataset branches are trained, Stage 2 requires only a few epochs to adjust. In our experiments, Stage 2 takes 40 minutes on two NVIDIA TITAN XP GPUs, while the baseline takes 5 hours. By default, we set the number of local regressors $G$ to 8, the overlap coefficient $\alpha$ to 2.0, the evidence learning regularization loss $\lambda$ to 0.01, the dataset split ratio in stages 1 and 2 to 4:1, and the starting MFF layer $K$ in the cross-dataset branch to 3.
@@
Table 1: Comparison of gaze estimation performance (angle error, °) in domain generalization. The best, second best, and third best results are denoted as red, blue, and green, respectively. † indicates the reproduced cross-domain gaze estimation results. Since the AGG code has not yet been open-sourced and certain results were not disclosed in their paper, we have left some sections of the tables unfilled.
Method Source Domain Target Domain Average
Type Name Train Data ETH G360 EyeDiap MPII Target All
Single-dataset Baseline Baseline ETH 4.55 20.70 10.70 8.46 9.58 11.10
Baseline G360 16.90 11.21 10.90 7.23 9.07 11.56
Cross-dataset Baseline Simple Mixing ETH+G360 4.68 11.42 6.42 6.95 6.69 7.37
Balanced Mixing ETH+G360 4.90 10.76 6.40 6.50 6.45 7.14
Multi-source Domain Generalization GMDG [32] ETH+G360 4.76 10.95 7.12 7.07 7.10 7.48
BNE [30] ETH+G360 5.44 11.06 6.73 6.51 6.62 7.44
Cross-domain Gaze Estimation PureGaze [11] ETH 4.77† 21.67† 7.48 7.08 7.28 10.25
PureGaze [11] G360 17.11† 11.21† 9.32 9.28 9.30 11.73
AGG [4] ETH 5.56 - 7.07 7.10 7.09 -
AGG [4] G360 - 13.03 7.93 7.87 7.90 -
Our Method EIF ETH+G360 4.42 10.46 6.45 6.01 6.23 6.84
4.3 Domain Generalization Results
We compared our method with four types of approaches: